# Scenario Analysis with Multivariate Bayesian Machine Learning Models

### v3

@@ 2.3. Generalized Impulse Response Functions
Lucas critique. Conditional forecasts and the GIRF variants relate to work on evaluating counterfactual policy rules, and may be exposed to the Lucas critique. This is the case particularly when imposing highly unusual conditioning scenarios, as opposed to mere “modest policy interventions,” which may leave behavioral patterns of economic agents unchanged (see Leeper and Zha, 2003). A recent discussion of these aspects is provided in McKay and Wolf (2023): the issue is that any imposed counterfactual paths on observables in this spirit require specific offsetting future shocks, and the restrictions hold only ex post and not necessarily in expectation. Since our approach is based on the same ideas as Antolin-Diaz et al. (2021), similar concerns as in linear models apply. It is worth noting that by virtue of a nonparametric approach, structural shocks may locally affect relationships among variables, even if the underlying functional form of the respective relationship remains fixed. Thus, while our approach is susceptible to the Lucas critique, it is likely more robust to related concerns than linear models.
@@ 3.1. Multivariate System Estimation
To estimate the model in (1) we rely on a conditional representation of its $n$ equations. Let $\bm{e}_{i}$ of size $1\times n$ denote the $i$th row of $\bm{I}_{n}$, and $\bm{E}_{i}$ of size $(n-1)\times n$ results from deleting the $i$th row of $\bm{I}_{n}$. Using $\bm{y}_{-it}=\bm{E}_{i}\bm{y}_{t}$ we may write $\bm{y}_{t}=\bm{e}_{i}^{\prime}y_{it}+\bm{E}_{i}^{\prime}\bm{y}_{-it}$. Under the assumptions of (1), $p(y_{it}\>|\>\bm{y}_{-it},\bullet)\propto\exp\left\{-(\bm{e}_{i}\bm{\Sigma}_{t}^{-1}\bm{e}_{i}^{\prime}y_{it}^{2}-2y_{it}\bm{e}_{i}\bm{\Sigma}_{t}^{-1}(\bm{F}(\bm{x}_{t})-\bm{E}_{i}^{\prime}\bm{y}_{-it}))/2\right\}$, which is a Gaussian with variance $\varsigma_{it}^{2}=(\bm{e}_{i}\bm{\Sigma}_{t}^{-1}\bm{e}_{i}^{\prime})^{-1}$ and mean ${\mu}_{it}=\varsigma_{it}^{2}(\bm{e}_{i}\bm{\Sigma}_{t}^{-1}(\bm{F}(\bm{x}_{t})-\bm{E}_{i}^{\prime}\bm{y}_{-it}))$. Indeed, this distribution is equivalent to a common representation of the conditional multivariate Gaussian (see, e.g., Cong et al., 2017, section 2). The mean can alternatively be written as ${\mu}_{it}=f_{i}(\bm{x}_{t})+\varsigma_{it}^{2}(\bm{e}_{i}\bm{\Sigma}_{t}^{-1}\bm{E}_{i}^{\prime})(\bm{y}_{-it}-\bm{E}_{i}\bm{F}(\bm{x}_{t}))$, and we define $\tilde{\mu}_{it}=\varsigma_{it}^{2}(\bm{e}_{i}\bm{\Sigma}_{t}^{-1}\bm{E}_{i}^{\prime})(\bm{y}_{-it}-\bm{E}_{i}\bm{F}(\bm{x}_{t}))$, i.e., $\mu_{it}=f_{i}(\bm{x}_{t})+\tilde{\mu}_{it}$. [^9] Related papers often either use a mapping between the structural and reduced form of the VAR to enable equation-by-equation estimation (see, e.g., Hauzenberger et al., 2025b), or rely on factor models for the reduced form errors (see, e.g., Clark et al., 2023). These approaches come with computational and inferential advantages and disadvantages. The former is simple to implement but requires parameterizing a structural form, which may cause issues such as inadvertently (instead of purposefully to achieve structural identification) breaking order-invariance of the equations. The latter allows for order-invariant inference but gives rise to the usual identification challenges of factor models. Our approach uses an order-invariant reduced form model for estimating the conditional mean relationships; see also Arias et al. (2023); Chan et al. (2024) for discussions. The $i$th equation of the multivariate model in regression form, conditional on all other equations, is then given by:
@@
which can be used in a Gibbs sampler to update the conditional mean relationships by looping through equations $i=1,\ldots,n$. This approach is similar to the one of Esser et al. (2024) and allows to treat each equation of the multivariate system individually, conditional on all other equations.
@@ 3.2. Bayesian Additive Regression Trees
The approach we discuss in Section 2 works with any implementation of multivariate models with additive and jointly Gaussian errors. That is, assuming a linear functional form for $\bm{F}(\bm{x}_{t})$ combined with suitable priors results in a standard Bayesian VAR (BVAR). [^10] When we consider linear versions of our model for comparisons, we implement this setting with $\bm{F}(\bm{x}_{t})=\bm{A}\bm{x}_{t}$ where $\bm{A}$ is an $n\times k$ matrix of reduced form VAR coefficients. We assume a horseshoe prior with a single global shrinkage component on these parameters, see also Hauzenberger et al. (2024); (12) can be used to update the VAR coefficients equation-by-equation from Gaussian posteriors. In case we treat $\bm{F}(\bm{x}_{t})$ nonparametrically, several options are available. Due to its versatility and established favorable empirical properties we mentioned earlier, we use BART to approximate the equation-specific functions in our applied work. That is, we consider a sum of $s=1,\ldots,S,$ tree functions $\ell_{is}(\bm{x}_{t}\>|\>\mathcal{T}_{is},\bm{\mathrm{m}}_{is})$ such that $f_{i}(\bm{x}_{t})\approx\sum_{s=1}^{S}\ell_{is}(\bm{x}_{t}\>|\>\mathcal{T}_{is},\bm{\mathrm{m}}_{is})$ where $\mathcal{T}_{is}$ are regression trees and $\bm{\mathrm{m}}_{is}$ is a vector of terminal node parameters (which serve as fitted values). Instead of having a single but complex tree, BART is akin to ensemble methods, and uses a sum of many simple trees (“weak learners”), which has been shown to work well.
@@ 3.3. Other Priors and Sampling Algorithm
Inspired by Carriero et al. (2024), we assume that $\bm{\Sigma}_{t}=s_{t}\bm{\Sigma}$ — the covariance structure only varies proportionally over time and $s_{t}$ is used to capture outliers. Our prior setup for the constant part of the covariance matrix follows Esser et al. (2024). Specifically, we use a hierarchical inverse Wishart prior $\bm{\Sigma}\>|\>\{a_{i}\}_{i=1}^{n}\sim\mathcal{W}^{-1}(s_{0},\bm{S}_{0}),$ where $s_{0}=\nu+n-1$, $\bm{S}_{0}=2\nu\cdot\text{diag}(1/a_{1},\ldots,1/a_{n})$ and $a_{i}\sim\mathcal{G}^{-1}(1/2,1/A_{j}^{2})$ for $i=1,\ldots,n,$ and a fixed scale parameter $A_{j}>0$. Setting $\nu=2$ implies a comparatively uninformative prior about the implied correlation structure, different from fixed-hyperparameter versions of this prior which has a tendency to overshrink. If applicable, we assume that $\Pr(s_{t}^{1/2}=1)=1-\mathfrak{p}$ and $\Pr(s_{t}^{1/2}\sim\mathcal{U}(2,\overline{\mathfrak{s}}))=\mathfrak{p}$, where $\mathcal{U}(2,\overline{\mathfrak{s}})$ is a discrete uniform distribution with (integer) support between $2$ and $\overline{\mathfrak{s}}=6$ and $\mathfrak{p}\sim\mathcal{B}(a_{\mathfrak{p}},b_{\mathfrak{p}})$ is the probability of observing an outlier. We set $a_{\mathfrak{p}}=1$, $b_{\mathfrak{p}}=50$ which a priori implies about 22% of the observations are outliers. Alternative and more flexible stochastic volatility specifications are available in this context (e.g., Chan, 2023). We obtain draws from the joint posterior of our model using a fairly straightforward MCMC algorithm. Details are provided in Appendix A.
@@ 4.2. Financial Conditions in the US and Tail Risk Scenarios
The scenarios are defined to reflect an increase of the NFCI by approximately $1$, 33 and $6$ unconditional standard deviations (SDs, reflecting tighter financial conditions) in 2018Q1, i.e., in $\mathcal{C}_{1}$, which we implement by placing these values as hard restriction on the NFCI in that quarter. From 2018Q2 onward we leave the future unrestricted, i.e., $\mathcal{C}_{h}=\varnothing$ for $h>1$. We investigate growth-at-risk (quantiles of real GDP), inflation-at-risk (quantiles of PCE inflation) and labor-at-risk (quantiles of growth in payroll employment) as our objects of interest (see Adams et al., 2021; Pfarrhofer, 2022; Botelho et al., 2024; Clark et al., 2024; Lopez-Salido and Loria, 2024, for related papers). The CF distributions (density estimates), for average growth rates $\overline{\bm{y}}_{T+h}=\sum_{j=1}^{h}\bm{y}_{T+j}/h$, are shown in Figure 3. In the lower panel, we show the difference between the conditional scenario distributions and the unconditional one, which yields the UGIRF (cumulated for all variables except the NFCI).
@@ A.2. Posterior Distributions and Sampling Algorithm
Updating all trees $s=1,\ldots,S,$ across equations $i=1,\ldots,n,$ yields an updated fit that can be used to compute the outlier-adjusted residuals $\bm{\epsilon}_{t}/\sqrt{s_{t}}=\bm{y}_{t}-\bm{F}(\bm{x}_{t})$. The posterior of the constant part of the covariance matrix is then given by:
$\bm{\Sigma}\>|\>\bullet\sim\mathcal{W}^{-1}\left(s_{0}+T,\bm{S}_{0}+\sum_{t=1}^{T}s_{t}^{-1}\bm{\epsilon}_{t}\bm{\epsilon}_{t}^{\prime}\right)$
@@
$\Pr(s_{t}^{1/2}=1\>|\>\bm{y}_{t},\bm{F}(\bm{x}_{t}),\bm{\Sigma},\mathfrak{p})$ $\propto\mathcal{N}(\bm{y}_{t}\>|\>\bm{F}(\bm{x}_{t}),\bm{\Sigma})\cdot(1-\mathfrak{p}),$
$\Pr(s_{t}^{1/2}=\mathfrak{s}\>|\>\bm{y}_{t},\bm{F}(\bm{x}_{t}),\bm{\Sigma},\mathfrak{p})$ $\propto\mathcal{N}(\bm{y}_{t}\>|\>\bm{F}(\bm{x}_{t}),\mathfrak{s}\bm{\Sigma})\cdot\left(\mathfrak{p}/(\overline{\mathfrak{s}}-1)\right),\quad\text{for }\mathfrak{s}=2,3,\ldots,\overline{\mathfrak{s}},$
@@ A.3. Recursive substitution and structural (G)IRFs
for iteration $m$ of our sampler. That is, we obtain a scaled version of the $j$th column of the impact matrix $\bm{H}^{-1}$ as in standard VAR models. We thus assume with our additive error specification that the impact of the respective shock is constant over time, and shocks of different signs and sizes are (proportionally) introduced by setting $d$ accordingly. This relates to scaled versions of the average causal effect for infinitesimally small shocks on impact, $\lim_{d\rightarrow 0}\bm{\delta}^{(d)}_{j\tau,1}/d=\bm{\beta}_{0}^{(m)}$, see also Kolesár and Plagborg-Møller (2025).
Next, using $\bm{\beta}_{0}^{(\texttt{s},m)}=\bm{H}^{-1}\bm{u}_{j\tau}^{(\texttt{s},m)}$ and $\bm{\beta}_{0}^{(\texttt{b},m)}=\bm{H}^{-1}\bm{u}_{j\tau}^{(\texttt{b},m)}$ to refer to the impact effects for the scenario and baseline forecasts, we can set $\bm{x}_{\tau+2}^{(\texttt{s},m)}=((\bm{y}_{\tau+1}+\bm{\beta}_{0}^{(\texttt{s},m)})^{\prime},\bm{y}_{\tau}^{\prime},\ldots,\bm{y}_{\tau-p+2}^{\prime})^{\prime}$ and $\bm{x}_{\tau+2}^{(\texttt{b},m)}=((\bm{y}_{\tau+1}+\bm{\beta}_{0}^{(\texttt{b},m)})^{\prime},\bm{y}_{\tau}^{\prime},\ldots,\bm{y}_{\tau-p+2}^{\prime})^{\prime}$. As in Section 2, using (5), we can iterate forward and obtain $\bm{\delta}_{\tau,2}^{(d,m)}=\bm{F}(\bm{x}_{\tau+2}^{(\texttt{s},m)})-\bm{F}(\bm{x}_{\tau+2}^{(\texttt{b},m)})$. [^1] Note that we may also compute other functions instead of the expectation, e.g., probabilities or quantiles, see also Gallant et al. (1993), and more recently, Jordà and Taylor (2025). For higher-order responses, iterate forward, draw from the conditional predictive distributions, and reconfigure $\bm{x}_{\tau+h}^{(\texttt{s},m)}$ and $\bm{x}_{\tau+h}^{(\texttt{b},m)}$ accordingly at each horizon. The structural GIRF is given by $\bm{\delta}_{\tau,h}^{(d,m)}=\bm{F}(\bm{x}_{\tau+h}^{(\texttt{s},m)})-\bm{F}(\bm{x}_{\tau+h}^{(\texttt{b},m)})$. This addresses future reduced form shocks to obtain the respective conditional expectation (implicitly via the recursive simulation of the scenario and baseline paths of the variables). Different to PGAS, this only requires a single realization instead of multiple particles.

### v4

@@ 2.3. Generalized Impulse Response Functions
Lucas critique. Conditional forecasts and the GIRF variants relate to work on evaluating counterfactual policy rules, and may be exposed to the Lucas critique. This is the case particularly when imposing highly unusual conditioning scenarios, as opposed to mere “modest policy interventions,” which may leave behavioral patterns of economic agents unchanged (see Leeper and Zha, 2003). A recent discussion of these aspects is also provided in McKay and Wolf (2023); a related issue is that any imposed counterfactual paths on observables in this spirit require specific neutralizing future shocks, and the restrictions hold only ex post and not necessarily in expectation. Since our framework is based on the idea of a future sequence of shocks delivering the scenario as in Antolin-Diaz et al. (2021), similar concerns as in linear models apply.
Two further remarks are relevant. First, by virtue of our nonparametric approach, structural shocks may locally affect relationships among variables, even if the underlying functional relationships remain constant over time. Thus, while our baseline setting is susceptible to the Lucas critique, it is likely more robust to related concerns than linear models. Second, Breitenlechner et al. (2024, see their Appendix C) recently recognized that the Lucas-robust approach of McKay and Wolf (2023) can be implemented in linear frameworks of the Antolin-Diaz et al. (2021) type by allowing offsetting shocks exclusively on impact. This restriction can straightforwardly be implemented in our setup, provided the matrix $\bm{H}$ is identified. In the spirit of our discussion in Section 2, to subsequently explore the future conditional moments of the observables stochastically, we further augmented the desired impact restriction with a restriction forcing all future structural shocks (for $h>1$) to adhere to their unconditional distribution.
@@ 3.1. Multivariate System Estimation
To estimate the model in (1) we rely on a conditional representation of its $n$ equations. Let $\bm{e}_{i}$ of size $1\times n$ denote the $i$th row of $\bm{I}_{n}$, and $\bm{E}_{i}$ of size $(n-1)\times n$ results from deleting the $i$th row of $\bm{I}_{n}$. Using $\bm{y}_{-it}=\bm{E}_{i}\bm{y}_{t}$ we may write $\bm{y}_{t}=\bm{e}_{i}^{\prime}y_{it}+\bm{E}_{i}^{\prime}\bm{y}_{-it}$. Under the assumptions of (1), $p(y_{it}\>|\>\bm{y}_{-it},\bullet)\propto\exp\left\{-(\bm{e}_{i}\bm{\Sigma}_{t}^{-1}\bm{e}_{i}^{\prime}y_{it}^{2}-2y_{it}\bm{e}_{i}\bm{\Sigma}_{t}^{-1}(\bm{F}(\bm{x}_{t})-\bm{E}_{i}^{\prime}\bm{y}_{-it}))/2\right\}$, which is a Gaussian with variance $\varsigma_{it}^{2}=(\bm{e}_{i}\bm{\Sigma}_{t}^{-1}\bm{e}_{i}^{\prime})^{-1}$ and mean ${\mu}_{it}=\varsigma_{it}^{2}(\bm{e}_{i}\bm{\Sigma}_{t}^{-1}(\bm{F}(\bm{x}_{t})-\bm{E}_{i}^{\prime}\bm{y}_{-it}))$. This distribution is equivalent to a common representation of the conditional multivariate Gaussian (see, e.g., Cong et al., 2017, Section 2). The mean can alternatively be written as ${\mu}_{it}=f_{i}(\bm{x}_{t})-\varsigma_{it}^{2}(\bm{e}_{i}\bm{\Sigma}_{t}^{-1}\bm{E}_{i}^{\prime})(\bm{y}_{-it}-\bm{E}_{i}\bm{F}(\bm{x}_{t}))$, and we define $\tilde{\mu}_{it}=-\varsigma_{it}^{2}(\bm{e}_{i}\bm{\Sigma}_{t}^{-1}\bm{E}_{i}^{\prime})(\bm{y}_{-it}-\bm{E}_{i}\bm{F}(\bm{x}_{t}))$, i.e., $\mu_{it}=f_{i}(\bm{x}_{t})+\tilde{\mu}_{it}$. [^9] Related papers often either use a mapping between the structural and reduced form of the VAR to enable equation-by-equation estimation (see, e.g., Hauzenberger et al., 2025b), or rely on factor models for the reduced form errors (see, e.g., Clark et al., 2023). These approaches come with computational and inferential advantages and disadvantages. The former is simple to implement but requires parameterizing a structural form, which may cause issues such as inadvertently (instead of purposefully to achieve structural identification) breaking order-invariance of the equations. The latter allows for order-invariant inference but gives rise to the usual identification challenges of factor models. Our approach uses an order-invariant reduced form model for estimating the conditional mean relationships (order-variant specifications may be consequential for conditional forecasts, see also Arias et al., 2023; Chan et al., 2024, for discussions). The $i$th equation of the multivariate model in regression form, conditional on all other equations, is then given by:
@@
which can be used in a Gibbs sampler to update the conditional mean relationships by looping through equations $i=1,\ldots,n$. This approach is similar to the one of Esser et al. (2024) and allows to treat each equation of the multivariate system individually, conditional on all other equations, which is computationally quick.
@@ 3.2. Bayesian Additive Regression Trees
The approach we discuss in Section 2 works with any implementation of multivariate models with additive and jointly Gaussian errors. That is, assuming a linear functional form for $\bm{F}(\bm{x}_{t})$ combined with suitable priors results in a standard Bayesian VAR (BVAR). [^10] When we consider linear versions of our model for comparisons, we implement this setting with $\bm{F}(\bm{x}_{t})=\bm{A}\bm{x}_{t}$ where $\bm{A}$ is an $n\times k$ matrix of reduced form VAR coefficients. We assume a horseshoe prior with a single global shrinkage component on these parameters, see also Hauzenberger et al. (2024); (12) can be used to update the VAR coefficients equation-by-equation from textbook Gaussian posteriors. In case we treat $\bm{F}(\bm{x}_{t})$ nonparametrically, several options are available. Due to its versatility and established favorable empirical properties we mentioned earlier, we use BART to approximate the equation-specific functions in our applied work. That is, we consider a sum of $s=1,\ldots,S,$ tree functions $\ell_{is}(\bm{x}_{t}\>|\>\mathcal{T}_{is},\bm{\mathrm{m}}_{is})$ such that $f_{i}(\bm{x}_{t})\approx\sum_{s=1}^{S}\ell_{is}(\bm{x}_{t}\>|\>\mathcal{T}_{is},\bm{\mathrm{m}}_{is})$ where $\mathcal{T}_{is}$ are regression trees and $\bm{\mathrm{m}}_{is}$ is a vector of terminal node parameters (which serve as fitted values). Instead of having a single but complex tree, BART is akin to ensemble methods, and uses a sum of many simple trees (“weak learners”), which has been shown to work well.
@@ 3.3. Other Priors and Sampling Algorithm
Inspired by Carriero et al. (2024), we assume that $\bm{\Sigma}_{t}=s_{t}^{2}\bm{\Sigma}$ — the covariance structure varies proportionally over time and $s_{t}$ is used to capture outliers. Our prior setup for the constant part of the covariance matrix follows Esser et al. (2024). Specifically, we use a hierarchical inverse Wishart prior $\bm{\Sigma}\>|\>\{a_{i}\}_{i=1}^{n}\sim\mathcal{W}^{-1}(s_{0},\bm{S}_{0}),$ where $s_{0}=\nu+n-1$, $\bm{S}_{0}=2\nu\cdot\text{diag}(1/a_{1},\ldots,1/a_{n})$ and $a_{i}\sim\mathcal{G}^{-1}(1/2,1/A_{j}^{2})$ for $i=1,\ldots,n,$ and a fixed scale parameter $A_{j}>0$. Setting $\nu=2$ implies a comparatively uninformative prior about the implied correlation structure, different from fixed-hyperparameter versions of this prior which has a tendency to overshrink. If applicable, we assume that $\Pr(s_{t}=1)=1-\mathfrak{p}$ and $\Pr(s_{t}\sim\mathcal{U}(2,\overline{\mathfrak{s}}))=\mathfrak{p}$, where $\mathcal{U}(2,\overline{\mathfrak{s}})$ is a discrete uniform distribution with (integer) support between $2$ and $\overline{\mathfrak{s}}=6$ and $\mathfrak{p}\sim\mathcal{B}(a_{\mathfrak{p}},b_{\mathfrak{p}})$ is the probability of observing an outlier. We set $a_{\mathfrak{p}}=1$, $b_{\mathfrak{p}}=50$ which a priori implies about 22% of the observations are outliers. Alternative and more flexible stochastic volatility specifications are available in this context (e.g., Chan, 2023). [^11] Another computationally attractive variant is to assume $\bm{\Sigma}_{t}=\bm{S}_{t}\bm{\Sigma}\bm{S}_{t}$ where $\bm{S}_{t}=\text{diag}(s_{1t},\ldots,s_{nt})$, which allows for variable-specific outlier detection using the same setup as in Section 3.3. These assumptions allow to invert $\bm{\Sigma}$ only once instead of having to invert $\bm{\Sigma}_{t}$ for all $t$ to compute the moments in (12). We obtain draws from the joint posterior of our model using a fairly straightforward MCMC algorithm. Details are provided in Appendix A.
@@ 4.2. Financial Conditions in the US and Tail Risk Scenarios
The scenarios are defined to reflect an increase of the NFCI by approximately $1$, 33 and $6$ unconditional standard deviations (SDs, reflecting tighter financial conditions) in 2018Q1, i.e., in $\mathcal{C}_{1}$, which we implement by placing these values as hard restriction on the NFCI in that quarter. From 2018Q2 onward we leave the future unrestricted, i.e., $\mathcal{C}_{h}=\varnothing$ for $h>1$. We investigate growth-at-risk (quantiles of real GDP), inflation-at-risk (quantiles of PCE inflation) and labor-at-risk (quantiles of growth in payroll employment) as our objects of interest (see Adams et al., 2021; Pfarrhofer, 2022; Clark et al., 2024; Lopez-Salido and Loria, 2024, for related papers). The CF distributions (density estimates), for average growth rates $\overline{\bm{y}}_{T+h}=\sum_{j=1}^{h}\bm{y}_{T+j}/h$, are shown in Figure 3. In the lower panel, we show the difference between the conditional scenario distributions and the unconditional one, which yields the UGIRF (cumulated for all variables except the NFCI).
@@ A.2. Posterior Distributions and Sampling Algorithm
Updating all trees $s=1,\ldots,S,$ across equations $i=1,\ldots,n,$ yields an updated fit that can be used to compute the outlier-adjusted residuals $\bm{\epsilon}_{t}/s_{t}=\bm{y}_{t}-\bm{F}(\bm{x}_{t})$. The posterior of the constant part of the covariance matrix is then given by:
$\bm{\Sigma}\>|\>\bullet\sim\mathcal{W}^{-1}\left(s_{0}+T,\bm{S}_{0}+\sum_{t=1}^{T}s_{t}^{-2}\bm{\epsilon}_{t}\bm{\epsilon}_{t}^{\prime}\right)$
@@
$\Pr(s_{t}=1\>|\>\bm{y}_{t},\bm{F}(\bm{x}_{t}),\bm{\Sigma},\mathfrak{p})$ $\propto\mathcal{N}(\bm{y}_{t}\>|\>\bm{F}(\bm{x}_{t}),\bm{\Sigma})\cdot(1-\mathfrak{p}),$
$\Pr(s_{t}=\mathfrak{s}\>|\>\bm{y}_{t},\bm{F}(\bm{x}_{t}),\bm{\Sigma},\mathfrak{p})$ $\propto\mathcal{N}(\bm{y}_{t}\>|\>\bm{F}(\bm{x}_{t}),\mathfrak{s}^{2}\bm{\Sigma})\cdot\left(\mathfrak{p}/(\overline{\mathfrak{s}}-1)\right),\quad\text{for }\mathfrak{s}=2,3,\ldots,\overline{\mathfrak{s}},$
@@ A.3. Recursive substitution and structural (G)IRFs
for iteration $m$ of our sampler. That is, we obtain a scaled version of the $j$th column of the impact matrix $\bm{H}^{-1}$ as in standard VAR models. We thus assume with our additive error specification that the impact of the respective shock is constant over time, and shocks of different signs and sizes are (proportionally) introduced by setting $d$ accordingly. This relates to scaled versions of the average effect for infinitesimally small shocks on impact, $\lim_{d\rightarrow 0}\bm{\delta}^{(d)}_{j\tau,1}/d=\bm{\beta}_{0}^{(m)}$, see also Kolesár and Plagborg-Møller (2025).
Next, using $\bm{\beta}_{0}^{(\texttt{s},m)}=\bm{H}^{-1}\bm{u}_{j\tau}^{(\texttt{s},m)}$ and $\bm{\beta}_{0}^{(\texttt{b},m)}=\bm{H}^{-1}\bm{u}_{j\tau}^{(\texttt{b},m)}$ to refer to the impact effects for the scenario and baseline forecasts, we can set $\bm{x}_{\tau+2}^{(\texttt{s},m)}=((\bm{y}_{\tau+1}+\bm{\beta}_{0}^{(\texttt{s},m)})^{\prime},\bm{y}_{\tau}^{\prime},\ldots,\bm{y}_{\tau-p+2}^{\prime})^{\prime}$ and $\bm{x}_{\tau+2}^{(\texttt{b},m)}=((\bm{y}_{\tau+1}+\bm{\beta}_{0}^{(\texttt{b},m)})^{\prime},\bm{y}_{\tau}^{\prime},\ldots,\bm{y}_{\tau-p+2}^{\prime})^{\prime}$. As in Section 2, using (5), we can iterate forward and obtain $\bm{\delta}_{\tau,2}^{(d,m)}=\bm{F}(\bm{x}_{\tau+2}^{(\texttt{s},m)})-\bm{F}(\bm{x}_{\tau+2}^{(\texttt{b},m)})$. [^1] Note that we may also compute other functions instead of the expectation, e.g., probabilities or quantiles, see also Gallant et al. (1993), and more recently, Jordà and Taylor (2025). For higher-order responses, iterate forward, draw from the conditional predictive distributions, and reconfigure $\bm{x}_{\tau+h}^{(\texttt{s},m)}$ and $\bm{x}_{\tau+h}^{(\texttt{b},m)}$ accordingly at each horizon. A draw from the structural GIRF distribution is given by $\bm{\delta}_{\tau,h}^{(d,m)}=\bm{F}(\bm{x}_{\tau+h}^{(\texttt{s},m)})-\bm{F}(\bm{x}_{\tau+h}^{(\texttt{b},m)})$. This addresses future reduced form shocks to obtain the respective conditional expectation (implicitly via the recursive simulation of the scenario and baseline paths of the variables). Different to PGAS, this only requires a single realization per MCMC iteration instead of multiple particles.